OpenStudy (anonymous):

Can someone just check my work? Find the derivitive of the given function: 2sin^-1(5x^4): I got 40x^3/sqrt(1-(5x^4)^2 tan^-1sqrt(5x) I got 5/(20x^2+2)sqrt(tan^-1*5x)

zepdrix (zepdrix):

Mmmm first one looks good. You applied chain rule. good good.

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}\arctan\left(\color{orangered}{\sqrt{5x}}\right)=\frac{1}{1+\left(\color{orangered}{\sqrt{5x}}\right)^2}\cdot \frac{1}{2\sqrt{5x}}\cdot 5\]Here is how the other problem would work out. First part is the arctan derivative. Then the second is from the chain rule, derivative of square root is 1 over 2 square roots, yes? Then the last part is from the 5x derivative.

OpenStudy (anonymous):

wow that's complicated so would it be 5/sqrtx *(10x+2) ?

zepdrix (zepdrix):

So you distributed the 2 and combined all the fractions? We should still have a sqrt5 somewhere in the denominator. Or you can simplify a little further since you have a 5 on top.\[\Large\rm \frac{\sqrt5}{\sqrt x(10x+2)}\]

OpenStudy (anonymous):

Thanks so much! do you mind looking over one more really quick thing? it's Determine the values of x for which the function is differentiable. y=sqrt(x^2+9) I got all values

OpenStudy (anonymous):

or is it -3,3,0?

zepdrix (zepdrix):

Ummm I think what we want to do is..

OpenStudy (anonymous):

first deriv and set = to 0?

zepdrix (zepdrix):

\[\Large\rm f(x)=\sqrt{x^2-9},\qquad\qquad\to\qquad\qquad f'(x)=\frac{x}{\sqrt{x^2-9}}\]Yah looks good, set the first derivative equal to zero. Then find out where the first derivative is `undefined`.

OpenStudy (anonymous):

it was +9

zepdrix (zepdrix):

Oh woops :\

OpenStudy (anonymous):

so all values?

zepdrix (zepdrix):

Mmmmmm yah I think so +_+ I don't see any troublesome spots.

OpenStudy (anonymous):

Thanks!